# Near-Discovery Observations of Interstellar Comet 3I/ATLAS with the NASA Infrared Telescope Facility

### v1

@@ 5.1 Ice and Dust
The dust distribution required to model the near-infrared spectrum of 3I and not get the visible reflectance completely wrong is anomalously steep and dominated by small grains compared to the comae of Solar System comets. The modeling challenges encountered might indicate something about 3I itself. It may be that 3I’s grains are somewhat different in composition than those of typical Solar System comets and thus amorphous carbon is a poor match to them. It is also possible that 3I is only weakly active, such that some of our reflectance spectrum is actually reflected light from the nucleus, and thus not modeled appropriately. If 3I really did brighten by half a magnitude or so around discovery (D. Z. Seligman et al., 2025), the dust around it at present may not be well described by a single power law. For 2I/Borisov, its dust grains might have been quite large, unlike many Solar System comets (B. Yang et al., 2021), so more complex spectral models will certainly be interesting to apply in the coming months. As 3I/ATLAS is observed at longer wavelengths, larger phase angles, and over a wider time span, these possibilities will become easier to assess.
@@ 5.2 Comparison to 2I/Borisov
Both of the (unambiguous) Interstellar Comets seem to lack the coma ice seen in many Solar System comets – particularly at the heliocentric distance at which we observed 3I. Our reflectance spectra showed no obvious signs of absorption features, and our spectral modeling constrained the areal fraction of water ice in the object’s coma to be approximately . If the grains in the coma were large and had mixed dust-and-ice compositions, somewhat more ice could be hidden in our data without showing any absorption features. Borisov (Q. Ye et al., 2020) and ATLAS have both shown cometary activity at distances in which water sublimation is inefficient, suggesting that their activity must be powered by the sublimation of a non-water substance like CO or CO2 (see, e.g., D. Bodewits et al. 2020; M. A. Cordiner et al. 2020 for Borisov). This would imply that they haven’t been warmed sufficiently to lose all of these more volatile species. These objects could still retain much of the water ice that they formed with – we just don’t see it in their comae. The distant activity and coma composition of comets is a critical way to begin to assess their thermal states and interior structures, and time will tell if the behavior and appearance of Borisov and ATLAS at large heliocentric distances are hints that Interstellar Comets store their ice differently than our own comets do. Until a more complete picture of 3I/ATLAS’s volatile inventory begins to form – as will likely come from observing it when it is at warmer temperatures near perihelion – it seems challenging to comment on the likelihood of M. J. Hopkins et al. (2025)’s suggestion that 3I may be intrinsically water rich based on its dynamics. If the object’s water ice were stored in large, mixed grains as opposed to the small pure grains proposed to be nearly-original in Solar System comets (J. M. Sunshine et al. 2007), its coma could potentially be water rich – this would just make it distinct from many Solar System comets.
@@ 6 Summary
3I/ATLAS’s lightcurve measured with ’Opihi on July 3 was low-amplitude and nearly consistent with flat. While the absence of light curve variability may indicate that 3I’snucleus has a roughly circular shape in projection (and thus a low lightcurve amplitude), it is also possible that 3I is surrounded by enough dust that any nuclear signal would be challenging to identify conclusively. Through ‘Opihi observations on July 4th, we measure a color of in the PANSTARRS system which is within of the red visible reflectivity measured by D. Z. Seligman et al. (2025) and C. Opitom et al. (2025), but is significantly redder than the colors measured by B. T. Bolin et al. (2025). Given 3I’s apparent activity, and non-D-type spectral features seen in the near-infrared, comparing 3I/ATLAS and solid surfaces of similar slopes (including 1I/’Oumuamua) may not be relevant.
Our NIR spectrum taken with SpeX also shows a red and linear reflectance spectrum like the D-type asteroids and comet nuclei, but only below 0.9 . At longer wavelengths, the spectrum becomes less red, then neutral (or possibly even slightly blue). No obvious absorption signatures, including those of water ice, are seen. The absence of absorption features is similar to 2I/Borisov (B. Yang et al., 2020), but the near-infrared spectra of the two interstellar comets show different NIR slopes and curvatures. The high amount of spectral curvature observed was difficult to reproduce with physically plausible spectral models, assuming it has grain compositions like Solar System comets and a single power law grain size distribution. Our most successful model retrieved an upper limit () on the areal fraction of water ice in the coma of 3I/ATLAS of , but combined dust and ice grains as opposed to physically separate populations of dust grains and ice grains may be able to push this number slightly higher. More work will be needed to identify 3I’s total volatile inventory to assess M. J. Hopkins et al. (2025)’s suggestion that it might be intrinsically water ice rich based on its dynamics. The distant activity of 2I and 3I implies they have some supervolatile ices left to sublimate, and thus that they have not presumably lost all of their less-volatile water ice. However, neither of the Interstellar Comets thus far have shown water ice in their comae despite clear activity beyond where ice sublimates easily, so perhaps these icy visitors store or release their ice(s) somewhat differently than many Solar System comets do. If

### v2

@@ 5.1 Ice and Dust
We found that the dust distribution required to model the near-infrared spectrum of 3I and not get the visible reflectance completely wrong is anomalously steep and dominated by small grains compared to the comae of Solar System comets. The modeling challenges encountered might indicate something about 3I itself. It may be that 3I’s grains are somewhat different in composition than those of typical Solar System comets and thus amorphous carbon is a poor match to them. Later observations at the IRTF reported in B. Yang et al. (2025) which showed similarly high amounts of spectral curvature were modeled with a combination of dust grains with compositions like that of the meteorite Tagish Lake and pure water ice, which would also be distinct from typical Solar System comet comae. It is also possible that 3I is only weakly active, such that some of our reflectance spectrum is actually reflected light from the nucleus, and thus not modeled appropriately. If 3I really did brighten by half a magnitude or so around discovery (D. Z. Seligman et al., 2025), the dust around it at present may not be well described by a single power law. For 2I/Borisov, its dust grains might have been quite large, unlike many Solar System comets (B. Yang et al., 2021), so more complex spectral models will certainly be interesting to apply in the coming months. As 3I/ATLAS is observed at longer wavelengths, larger phase angles, and over a wider time span, these possibilities will become easier to assess.
@@ 5.2 Comparison to 2I/Borisov
Both of the (unambiguous) Interstellar Comets seem to lack the coma ice seen in many Solar System comets – particularly at the heliocentric distance at which we observed 3I. Our reflectance spectra showed no obvious signs of absorption features, and our spectral modeling constrained the areal fraction of water ice in the object’s coma to be approximately . If the grains in the coma were large and had mixed dust-and-ice compositions, somewhat more ice could be hidden in our data without showing any absorption features. Observations taken approximately ten days after ours were reported in B. Yang et al. (2025) to show some evidence of water ice, but more work will be needed to disentangle how much is variabilty in the comet on short timescales, how much is due to changes in the comet’s activity as it approached the sun, and how much is due to differences in our modeling approaches. Borisov (Q. Ye et al., 2020) and ATLAS have both shown cometary activity at distances in which water sublimation is inefficient, suggesting that their activity must be powered by the sublimation of a non-water substance like CO or CO2 (see, e.g., D. Bodewits et al. 2020; M. A. Cordiner et al. 2020 for Borisov). This would imply that they haven’t been warmed sufficiently to lose all of these more volatile species. These objects could still retain much of the water ice that they formed with – we just don’t see it in their comae. The distant activity and coma composition of comets is a critical way to begin to assess their thermal states and interior structures, and time will tell if the behavior and appearance of Borisov and ATLAS at large heliocentric distances are hints that Interstellar Comets store their ice differently than our own comets do. Until a more complete picture of 3I/ATLAS’s volatile inventory begins to form – as will likely come from observing it when it is at warmer temperatures near perihelion – it seems challenging to comment on the likelihood of M. J. Hopkins et al. (2025)’s suggestion that 3I may be intrinsically water rich based on its dynamics. If the object’s water ice were stored in large, mixed grains as opposed to the small pure grains proposed to be nearly-original in Solar System comets (J. M. Sunshine et al. 2007), its coma could potentially be water rich – this would just make it distinct from many Solar System comets.
@@ 6 Summary
3I/ATLAS’s lightcurve measured with ’Opihi on July 3 was low-amplitude and nearly consistent with flat. While the absence of light curve variability may indicate that 3I’s nucleus has a roughly circular shape in projection (and thus a low lightcurve amplitude), it is also possible that 3I is surrounded by enough dust that any nuclear signal would be challenging to identify conclusively. That said, longer rotation periods like those suggested in R. de la Fuente Marcos et al. (2025) might be reconcilable with our data. Through ‘Opihi observations on July 4, we measure an averaged color of which is within of the red visible reflectivity measured by D. Z. Seligman et al. (2025); C. Opitom et al. (2025); R. de la Fuente Marcos et al. (2025); B. T. Bolin et al. (2025); M. Belyakov et al. (2025). Given 3I’s apparent activity, and non-D-type spectral features seen in the near-infrared, comparing 3I/ATLAS and solid surfaces of similar slopes (including 1I/’Oumuamua) may not be relevant.
Our NIR spectrum taken with SpeX also shows a red and linear reflectance spectrum like the D-type asteroids and comet nuclei, but only below 0.9 . At longer wavelengths, the spectrum becomes less red, then neutral (or possibly even slightly blue). No obvious absorption signatures, including those of water ice, are seen. The absence of absorption features is similar to 2I/Borisov (B. Yang et al., 2020), but the near-infrared spectra of the two interstellar comets show different NIR slopes and curvatures. The high amount of spectral curvature observed was difficult to reproduce with physically plausible spectral models, assuming it has grain compositions like Solar System comets and a single power law grain size distribution. Our most successful model retrieved an upper limit () on the areal fraction of water ice in the coma of 3I/ATLAS of , but combined dust and ice grains as opposed to physically separate populations of dust grains and ice grains may be able to push this number slightly higher. More work will be needed to identify 3I’s total volatile inventory to assess M. J. Hopkins et al. (2025)’s suggestion that it might be intrinsically water ice rich based on its dynamics, especially in light of B. Yang et al. (2025)’s later potential detection of icy grains in 3I/ATLAS when it was brighter and somewhat more active. The distant activity of 2I and 3I implies they have some supervolatile ices left to sublimate, and thus that they have not presumably lost all of their less-volatile water ice. However, neither of the Interstellar Comets thus far have shown water ice in their comae despite clear activity beyond where ice sublimates easily, so perhaps these icy visitors store or release their ice(s) somewhat differently than many Solar System comets do.